Add. Express your result in simplest form.
4/xy + 6/xy
My answer is: 10/ x^2 y^2 am I right?


Answer by stanbon(48568)

No. When the denominators are the same adding or subtracting
fractions is like adding or subtracting cars.
4/xy + 6/xy = 10/xy
just like
4 cars + 6 cars = 10 cars.
---------
Similarly
6/xy - 4/xy = 2/xy
just like
6cars - 4cars = 2cars
===========
cheers,
Stan H.